Professor Grouch has developed an abundance index based on Snuffle poop abundance (once a grouch always...

  1. Professor Grouch has developed an abundance index based on Snuffle poop abundance (once a grouch always a grouch..). Here is a calibration model he developed that correlates snuffle poop density (PSF) to actual snuffle density (NSF) (units are number/ha):    PSF = 10*NSF+5. Based on transects Telly estimates the average poop density to be 20 poops/ha. What is his estimate of Snuffle density (snuffles/ha)?
    1. 0.02
    2. 0.75
    3. 1.5
    4. 15
    5. 20
Answer #1

Answer : c)1.5

In the given problem, it is mentioned that

PSF= 10* NSF+5

where

PSF is snuffle poop density and NSF is actual snuffle density.

It is estimated that the average poop density(PSF) is 20 poops/ha. The given model of correlation can be used to determine the snuffle density(NSF).

20 = 10 * NSF + 5 .....1

by rearranging equation 1 we get ,

=> NSF = (20-5)/10

NSF = 1.5 snuffles/ha